OpenStudy (anonymous):

Algebra II Question drawn below. Can anyone double-check my answer of x^13/30.

OpenStudy (anonymous):



OpenStudy (anonymous):

yes, \[\frac{5}{6}-\frac{2}{5}=\frac{13}{30}\]

OpenStudy (anonymous):

k thx :)